OpenStudy (anonymous):

We are studying elastic collisions. There is this system of equations problem where we are solving for the first and second final velocities. Help! equations: m1v1,i+m2v2i = m1v1f+m2v2f 0.5m1v1i^2+0.5m2v2i^2 = 0.5m1v1f^2 + 0.5m1v1f^2 Givens: m1= 2kg m2=3kg v1,i= 4m/s v2,i=-5m/s unknown: v1,f= ? v2,f=? Please help!

OpenStudy (perl):

we can plug in

OpenStudy (anonymous):

Then I simplified one side and used the substitution technique

OpenStudy (perl):

yes that looks correct

OpenStudy (perl):

what you did, you might have solved for v_2_f

OpenStudy (anonymous):

I have the same expression but the 2 and 3 are swapped. Can you please solve this problem?

OpenStudy (perl):

$$ \Large { m_1 v_{1_i} + m_2 v_{2_i} = m_1 v_{1_f} + m_2 v_{2_f} \\ 0.5~m_1v_{1_i}^{2}+0.5~m_2v_{2_i}^2 = 0.5~m_1v_{1_f}^2+0.5~m_2v_{2_f}^2 \\ \therefore \\ 2(4) + 3(-5) = 2( v_{1_f} ) + 3(v_{2_f}) \\0.5 (2)(4^{2}) + 0.5 (3) (-5)^{2} = 0.5 (2) v_{1_f}^{2} + 0.5(3) v_{2_f}^{2} \\ \therefore \\ -7 = 2( v_{1_f} ) + 3(v_{2_f}) \\53.5 = 1~ v_{1_f}^{2} + 2.5~ v_{2_f}^{2} \\ \therefore \\ -7 - 3(v_{2_f}) = 2( v_{1_f} )\\ \frac{-7 - 3(v_{2_f})}{2} = v_{1_f} \\ \therefore \\ 53.5 =1\cdot \left ( \frac{-7 - 3(v_{2_f})}{2} \right)^2 + 2.5~ v_{2_f}^{2} } $$

OpenStudy (perl):

why are the 2 and 3 swapped?

OpenStudy (anonymous):

I probably plugged them in incorrectly.

OpenStudy (perl):

m1 = 2, m2 = 3 . i might have made a mistake as well, so double check that

OpenStudy (perl):

m1 = 2kg , m2 = 3kg v_1_i = 4 m/s , v_2_i = -5 m/s

OpenStudy (perl):

ok so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

where did the four in the last line come from?

OpenStudy (perl):

$$ \Large { m_1 v_{1_i} + m_2 v_{2_i} = m_1 v_{1_f} + m_2 v_{2_f} \\ 0.5~m_1v_{1_i}^{2}+0.5~m_2v_{2_i}^2 = 0.5~m_1v_{1_f}^2+0.5~m_2v_{2_f}^2 \\ \therefore \\ 2(4) + 3(-5) = 2( v_{1_f} ) + 3(v_{2_f}) \\0.5 (2)(4^{2}) + 0.5 (3) (-5)^{2} = 0.5 (2) v_{1_f}^{2} + 0.5(3) v_{2_f}^{2} \\ \therefore \\ -7 = 2( v_{1_f} ) + 3(v_{2_f}) \\53.5 = 1~ v_{1_f}^{2} + 2.5~ v_{2_f}^{2} \\ \therefore \\ -7 - 3(v_{2_f}) = 2( v_{1_f} )\\ \frac{-7 - 3(v_{2_f})}{2} = v_{1_f} \\ \therefore \\ 53.5 =1\cdot \left ( \frac{-7 - 3(v_{2_f})}{2} \right)^2 + 2.5~ v_{2_f}^{2} \\ \therefore \\ 53.5 =\frac{\left ( -7 - 3(v_{2_f})\right)^2}{2^2} + 2.5~ v_{2_f}^{2} \\ \therefore \\ \color{red} {2^2} \cdot 53.5 = \color{red} {2^2}\left( \frac{\left ( -7 - 3(v_{2_f})\right)^2}{2^2} + 2.5~ v_{2_f}^{2} \right ) \\ \therefore \\ 2^2\cdot 53.5 =\left ( -7 - 3(v_{2_f})\right)^2 + 2^2\cdot 2.5~ v_{2_f}^{2}\\ \\ \therefore \\ 4\cdot 53.5 = (-7)^2 + 2(-7)(-3)v_{2_f} + 9v_{2_f}^2 + 4\cdot 2.5~ v_{2_f}^{2} } $$

OpenStudy (anonymous):

okay I see

OpenStudy (anonymous):

lets solve this and then use the link to verify if it works

OpenStudy (anonymous):

never mind

OpenStudy (perl):

Also this equation is found here http://en.wikipedia.org/wiki/Elastic_collision#One-dimensional_Newtonian it is easier to label u1 , u2 for initial speeds. v1,v2 for final speeds

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

Vi= -6.8m/s I think

OpenStudy (anonymous):

yes! wolfram alpha had similar results let me check the other value

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is the link ok

OpenStudy (perl):

sorry i made a typo when i put in the equation. its fixed now

OpenStudy (anonymous):

It is totally fine. I got 2.4 m/s for v2 but the wolfram alpha calculation says 3.7

OpenStudy (perl):

so let me go back and simplify this equation, see if we can solve

OpenStudy (perl):

$$\Large { 2\cdot 4 + 3\cdot (-5) = 2x + 3y \\ \frac{1}{2} \cdot 2\cdot 4^2 + \frac12\cdot 3(-5)^2 = \frac12 2 \cdot x^2 + \frac12 3 \cdot y^2 } $$

OpenStudy (perl):

multiply both sides of that second equation by 2, to remove the 1/2

OpenStudy (perl):

$$ \Large { 2\cdot 4 + 3\cdot (-5) = 2x + 3y \\ \frac{1}{2} \cdot 2\cdot 4^2 + \frac12\cdot 3(-5)^2 = \frac12 2 \cdot x^2 + \frac12 3 \cdot y^2 \\ \iff \\ 2\cdot 4 + 3\cdot (-5) = 2x + 3y \\ 2\cdot 4^2 + 3(-5)^2 = 2 \cdot x^2 + 3 \cdot y^2 \\ \iff \\ -7 = 2x + 3y\\ 107 = 2x^2 + 3y^2 \\ \therefore \\ 107 = 2x^2 + 3 \left( \frac{-7-2x}{3}\right )^2 \\ \iff \\ 107 = 2x^2 + 3 \cdot \frac{(7+2x)^2}{3^2} \\ \\107 = 2x^2 + \frac{(7+2x)^2}{3} \\ \\ \iff \text {multiply both sides by 3 } \\ \\3\cdot 107 = 3\cdot 2x^2 + (7+2x)^2\\ \iff \\ \\ 321 = 6x^2 + 7^2 + 28x + 4x^2 \\ \iff \\ \\ 321 = 10x^2 + 28x + 49 \\ \iff \\ \\ 0 = 10x^2 + 28x -272 \\ \iff \\ \\ 0 = 5x^2 + 14x -136 \\ \iff \\ 0 = ( 5x + 34) (x - 4) \\ \iff \\ x = -34/5 , 4 } $$

OpenStudy (perl):

It is better to solve the problem generally and use the formula derived, than to solve this equation for specific values. also you can reject the solution x = 4, y = -5 since that implies there was no collision

OpenStudy (error1603):

right